Mrs. Anderson gave an exam to her math class, the exam was worth 225 points, and it contained 50 questions. Some of the questions were worth 3 points, and some questions were worth 8 points. How many of each question were on the exam?

Answer by Boreal(15235)

x=3 point questions
y=8 point questions
x+y=50
3x+8y=225
----------
multiply the top by -3
-3x-3y= -150
3x+8y=225
5y=75
y=15
x=35
105 points from x
120 points from y
They add to 225.